# Guaranteed Recovery of Unambiguous Clusters

### v1

@@ Abstract
Clustering is often a challenging problem because of the inherent ambiguity in what the “correct” clustering should be. Even when the number of clusters $K$ is known, this ambiguity often still exists, particularly when there is variation in density among different clusters, and clusters have multiple relatively separated regions of high density. In this paper we propose an information-theoretic characterization of when a $K$-clustering is ambiguous, and design an algorithm that recovers the clustering whenever it is unambiguous. This characterization formalizes the situation when two high density regions within a cluster are separable enough that they look more like two distinct clusters than two truly distinct clusters in the clustering. The algorithm first identifies $K$ partial clusters (or “seeds”) using a density-based approach, and then adds unclustered points to the initial $K$ partial clusters in a greedy manner to form a complete clustering. We implement and test a version of the algorithm that is modified to effectively handle overlapping clusters, and observe that it requires little parameter selection and displays improved performance on many datasets compared to widely used algorithms for non-convex cluster recovery.
@@ I Introduction
In this paper, we try to overcome these difficulties from an information theoretic perspective by proposing a necessary condition for a $K$-clustering $C$ to be unambiguous, along with an algorithm that recovers $C$ whenever this condition holds. This condition characterizes the situation where two high density regions within the same cluster in $C$ look more like two distinct clusters than two truly distinct clusters in $C$. Our framework yields the provable recovery of unambiguous $K$-clusterings that can have
@@
Our approach is information theoretic in the sense that our algorithm guarantees recovery of a $K$-clustering whenever the dataset warrants an unambiguous $K$-clustering. This is in contrast to algorithms like $K$-means that are motivated by the optimization of a clustering quality measure.
The clustering algorithm we propose is somewhat complex, but has a computationally efficient runtime. It works by first finding a small subset of points called a “seed” from each cluster, and then expanding the seeds to form clusters. We implement and test a version of the algorithm that is modified to handle overlapping clusters well, and observe that it requires little parameter selection, and delivers improved performance on artificial and benchmark datasets compared to widely used algorithms for non-convex cluster recovery.
@@ II Related Work
Our algorithm works by first identifying $K$ partial clusters (or “seeds”) using a density-based approach, and then adding unclustered points to the initial $K$ partial clusters in a greedy manner to form a complete clustering. To extract these seeds, we sequentially find seeds of decreasing density. The intuitive idea of sequentially finding disjoint clusters of decreasing density has been employed in various work [11, 12]. However, to the best of our knowledge, our algorithm for finding $K$ seeds and our algorithm for expanding the seeds to form complete clusters have not previously appeared in the literature. Furthermore, we are not aware of any existing mathematical analysis of density-based clustering algorithms that give guarantees similar to those presented in this paper.
@@ III Preliminaries
We say a point $x_{1}$ is $\epsilon$-connected to a point $x_{t}$ if there exists some sequence of points $x_{1},x_{2},...,x_{t}$ such that $x_{i+1}$ is distance at most $\epsilon$ from $x_{i}$ for $11<i<t-$ 1 and $\epsilon_{N_{p}}(x_{i})\leq\epsilon$ for $1<i<t$. A set of points $c$ is called $\epsilon$-connected if every pair of points in $c$ is $\epsilon$-connected.
For a given $\epsilon$, a set of points $c$ is called an $\epsilon$-cluster if it is a maximal $\epsilon$-connected set of points. In other words, any point that is $\epsilon$-connected to a point in an $\epsilon$-cluster $c$ is included in $c$. A set of points $c$ is called a maximal cluster if it is an $\epsilon$-cluster for some $\epsilon$. For a given $\epsilon$, it is helpful to consider the graph that is formed where each node corresponds to a data point, and an edge is drawn between two nodes if the corresponding points $x_{1},x_{2}$ are such that $\epsilon_{N_{p}}(x_{1})\leq\epsilon$, $\epsilon_{N_{p}}(x_{2})\leq\epsilon$, and $d(x_{1},x_{2})\leq\epsilon$. A set of points is an $\epsilon$-cluster if and only if it corresponds to a connected component in this graph.
@@
For a node $v\in V$, $c(v)$ denotes the cluster corresponding to $v$. Each node $v$ has a real number $\epsilon(v)$ associated with it where $\epsilon(v)$ is the smallest $\epsilon$ such that $c(v)$ is a $\epsilon$-cluster. For a leaf node $v$ corresponding to the cluster $\{x\}$, $\epsilon(v)=\epsilon_{N_{p}}(x)$. A node $v_{0}$ has children $v_{1},v_{2},...,v_{i}$ if $c(v_{1}),c(v_{2}),...,c(v_{i})$ form the smallest possible partition of $c(v_{0})$ composed of maximal clusters, which is guaranteed to be unique. The root node $v_{r}$ of the resulting tree is such that $c(v_{r})=X$.
@@ IV-A Weak Separability
For a given $N_{p}$, $C$ is called weakly separable if for each $c\in C$, $c$ is $\epsilon$-connected where $\epsilon<\min_{c^{\prime}\in C,\;c^{\prime}\neq c}\epsilon(c,c^{\prime})$.
If a clustering $C$ is not weakly separable, there is ambiguity in what the correct clustering should be in the sense that there must exist a cluster $c\in C$ such that any $\epsilon$-cluster containing $c$ must include at least one point from a distinct cluster $c^{\prime}\in C$. Figures 1 and 2 illustrate weak separability.
@@
If there are two weakly separable clusterings for $X$, it is impossible to guarantee that we recover one of them if our only criteria is to find a weakly separable clustering. Because there usually does not exist a unique weakly separable $K$-clustering even when an intuitively correct $K$-clustering exists, weak separability alone as a sufficient condition for clustering recoverability is not adequate.
@@ IV-B Local Maximum Separability
LM-separability specifies that for every local maximum $x\in X$ where there exists a $y\in c(x)$ whose density is at least as high as that of $x$, the minimum relative separability of $x$ to such a $y$ is smaller than the relative separability of the highest density point in any cluster to another cluster. This is a condition that should hold for a clustering to unambiguous because if there exists a local maximum $x$ where the minimum relative separability to such a point $y\in c(x)$ is higher than the relative separability of a highest density point $z\in X_{c}^{*}$ for some $c\in C$ to another $c^{\prime}\in C$ (and therefore to a local maximum in $c^{\prime}$), then in a sense, $x$ looks more like it belongs to a separate cluster from $y$ than $z$ looks like it belongs to a separate cluster from $c^{\prime}$, and the correct $K$-clustering of $X$ is ambiguous. Figures 1 and 2 illustrate LM-separability.
Figure 1: A plot showing the sparsity value $\epsilon_{3}(x)$ of each point $x\in X$ in a 60 point one-dimensional clustering $C=[c_{1},c_{2}]$ that is weakly separable and LM-separable for $N_{p}=3$, implying that recovery is guaranteed. The orange, red, and pink stars correspond to local maximums. The red stars correspond to points in $X_{c}^{*}$ for $c\in C$. The pink star corresponds to the local maximum $x\in X$ such that $\min_{y\in c(x):\;\epsilon_{N_{p}}(y)\leq\epsilon_{N_{p}}(x)}A(x,y)=A^{\ell}(C)$. $C$ is weakly separable because $\epsilon^{*}(c_{1})=0.74$ and $\epsilon^{*}(c_{1})=1.14$, while $\min_{c,c^{\prime}\in C}\epsilon(c,c^{\prime})=1.18$. $C$ is LM-separable because $A^{\ell}(C)=8.95$ while $\min_{c,c^{\prime}\in C}\min_{z\in X_{c}^{*}}A(z,c^{\prime})=15.35$.
Note that LM-separability does not imply weak separability. Let $N_{p}=2$, and consider the clustering $C=[\{7,8,10,13,21\},\{17,25,27\}]$. $C$ is clearly not weakly separable because the clusters overlap, but is LM-separable, as $7,8,25,27$ are the only local maximums.
@@
For a given $N_{p}$, it is in general not possible to recover the unique weakly separable, LM-separable clustering by simply finding an extension of the (partial) clustering given by the $\epsilon$-cut of the dendrogram that contains $K$ clusters (if there exists such an $\epsilon$) as proved in Lemma 2. Since DBSCAN follows this approach, it is not sufficient for finding the unique weakly separable, LM separable clustering.
@@ V Algorithm and Proof of Theorem 1
Algorithm 1 Minimal Seed Expansion
@@
1:  $C^{\prime}\!\leftarrow$ output of Algorithm 2 with parameters $N_{p},M,D,K$
2:  $\hat{C}\leftarrow$ output of Algorithm 3 with parameters $X,C^{\prime},N_{p}$
@@
The min-$A$ clustering of $X$ with parameters $N_{p},M,D,K$ is given by $C_{g}(X,N_{p},A_{min},M,D)$ where
$A_{min}=\min_{A\text{ such that }|C_{g}(X,N_{p},A,M,D)|=K}A.$
For parameters $N_{p},M,D,K$, the min-$A$ clustering is denoted by $C_{m}(X,N_{p},M,D,K)$. $A_{min}\geq 1$ since $c^{*}(x,\epsilon)=\{\}$ if $\epsilon<\epsilon_{N_{p}}(x)$.
Algorithm 1 accepts $N_{p},M,D$ along with a number of clusters $K$ as input, and begins by finding the min-$A$ partial clustering with parameters $M,D,K$. The min-$A$ partial clustering is then passed to Algorithm 3 to produce the final output clustering.
@@
        $\text{Tried}\leftarrow\text{Tried}\cup x$
@@
  $Y\leftarrow X\setminus(\cup_{c\in C^{\prime}})$
@@
For a given $N_{p},$ if $C$ is LM-separable, then $|C_{g}(X,N_{p},A^{\ell}(C),1,\infty)|=K$.
@@
We will prove that all partial clusters output by Algorithm 2 are subsets of distinct clusters in $C$. Consider the $i$th partial cluster output by Algorithm 2. The highest density point $x$ that the algorithm uses to build the $i$th cluster is clearly a local maximum, and can either be from a previously partially reconstructed cluster in $C$, or can be from a new cluster in $C$. If $x$ is from a previously partially reconstructed cluster $c^{\prime}\in C$, then it must include some $z\in X_{c^{\prime}}^{*}$ by the definitions of LM-separability and $A^{\ell}(C)$, and the fact that $\epsilon_{N_{p}}(x)\geq\epsilon_{N_{p}}(z).$ This is a contradiction. This along with the fact that $M=1\leq N_{p}$ guarantees that $x$ must be from a new cluster $c^{\prime\prime}\in C$ and $x\in X_{c^{\prime\prime}}^{*}$. The $i$th partial cluster $c^{*}(x,A^{\ell}(C)\cdot\epsilon_{N_{p}}(x))$ must only include points from $c^{\prime\prime}$ by the definition of LM-separated and $A^{\ell}(C)$. Because only a new cluster center can be added to the partial clustering in each iteration of the algorithm, a partial $K$-clustering is output. ∎
@@
Therefore, the algorithm outputs non-overlapping partial clusters such that each partial cluster chosen does not include any points from true clusters that are not yet part of a previously output partial cluster. Observe that, if no points in a cluster $c$ have been added to previously output partial clusters by the time $w\in X_{c}^{*}$ is selected as the highest density unclustered point by the algorithm, then a partial cluster centered at $w$ that is a non-empty subset of $c$ will be output by the algorithm in that step since $M=1\leq N_{p}$, $A_{min}\leq A^{\ell}(C)$ and $C$ is LM-separable. Therefore, we obtain a partial clustering that contains a partial cluster centered at $w\in X_{c}^{*}$ that is a non-empty subset of $c$ for each $c\in C$. This is a min-$A$ partial clustering for $K$, so it is a partial $K$-clustering that contains a partial cluster centered at some $w\in X_{c}^{*}$ that is a non-empty subset of $c$ for each $c\in C$. Thus, the algorithm outputs a partial $K$-clustering that is extendable to $C$. ∎
@@
For a given $X,N_{p},M,D,$ the number of clusters in $C_{g}(X,N_{p},A,M,D)$ monotonically decreases as $A$ increases because of the following property. Consider some $a,a^{\prime}\geq 1$ such that $a>a^{\prime}$. At the end of the $i$th iteration of Algorithm 2, the set of remaining candidates for future $x^{*}$ where $c^{*}(x^{*},A\cdot\epsilon_{N_{p}}(x^{*}))$ will not intersection with a previously output cluster in the case when $A=a^{\prime}$ is a superset of the set of such candidates when $A=a$.
@@
Observe that for each greedy step of Algorithm 3, any point $x\in X\setminus(\cup_{c\in\hat{C}}c)$ and cluster $c\in\hat{C}$ that minimize the quantity $\epsilon^{!}(x,c)=\min_{y\in c}\max(d(x,y),\epsilon_{N_{p}}(x),\epsilon_{N_{p}}(y))$ can be selected as $(x^{*},c^{*})$. To see this, consider the set $S$ of unclustered points that are closest to some cluster in terms of $\epsilon$-distance. Denote this minimum $\epsilon$-distance by $\epsilon^{\prime}$. Clearly, $S$ must include an unclustered point that contains a clustered point in its $\epsilon^{\prime}$-ball. Thus, at each greedy step of Algorithm 3, we can pick $(x,c)$ that minimizes $\epsilon^{!}(x,y)$.
@@
For each greedy step of Algorithm 3, we simply set $(x^{*},c^{*})$ equal to the $(x,c)$ that minimizes $\arg\min_{c\in\hat{C}}\epsilon^{!}(x,c)$ in $O(|X|)$ time by checking $(\epsilon^{!}(x)$, $c^{!}(x))$ for all $x\in X\setminus(\cup_{c\in\hat{C}}c)$. After assigning $x^{*}$ to $c^{*}$, we update $\epsilon^{!}(x)$ for every remaining unclustered point $x$ by setting $\epsilon^{!}(x)\leftarrow\min(\epsilon^{!}(x),\;\max(d(x,x^{*}),\;\epsilon_{N_{p}}(x),\;\epsilon_{N_{p}}(x^{*})))$, and setting $c^{!}(x)\leftarrow c^{*}$ if the value of $\epsilon^{!}(x)$ is changed. Each of these greedy steps takes $O(|X|)$ time and there are at most $O(|X|)$ greedy steps. This stage of the algorithm therefore runs in $O(|X|^{2})$ time. ∎

### v3

@@ Abstract
Clustering is often a challenging problem because of the inherent ambiguity in what the “correct” clustering should be. Even when the number of clusters $K$ is known, this ambiguity often still exists, particularly when there is variation in density among different clusters, and clusters have multiple relatively separated regions of high density. In this paper we propose an information-theoretic characterization of when a $K$-clustering is ambiguous, and design an algorithm that recovers the clustering whenever it is unambiguous. This characterization formalizes the situation when two high density regions within a cluster are separable enough that they look more like two distinct clusters than two truly distinct clusters in the $K$-clustering. The algorithm first identifies $K$ partial clusters (or “seeds”) using a density-based approach, and then adds unclustered points to the initial $K$ partial clusters in a greedy manner to form a complete clustering. We implement and test a version of the algorithm that is modified to effectively handle overlapping clusters, and observe that it requires little parameter selection and displays improved performance on many datasets compared to widely used algorithms for non-convex cluster recovery.
@@ I Introduction
In this work, we try to overcome these difficulties from an information-theoretic perspective by proposing a condition that should hold for a $K$-clustering $C$ to be considered unambiguous, along with an algorithm that recovers $C$ whenever this condition holds. This condition characterizes the situation where two high density regions within the same cluster in $C$ look more like two distinct clusters than two truly distinct clusters in $C$. Our framework yields the provable recovery of unambiguous $K$-clusterings that can have
@@
Our approach is information-theoretic in the sense that our algorithm guarantees recovery of a $K$-clustering whenever the dataset warrants an unambiguous $K$-clustering in our framework. This is in contrast to algorithms like $K$-means that are motivated by the optimization of a clustering quality measure.
The clustering algorithm we propose is somewhat complex, but has a computationally efficient runtime. It works by first finding a small subset of points called a “seed” from each cluster, and then expanding the seeds to form clusters. We implement and test a version of the algorithm that is modified to handle overlapping clusters well, and observe that it requires little parameter selection, and delivers improved performance on artificial and benchmark datasets compared to widely used algorithms for non-convex cluster recovery. This implementation is available at: https://github.com/kmazooji/Minimal-Seed-Expansion.
@@ II Related Work
Our algorithm works by first identifying $K$ partial clusters (or “seeds”) using a density-based approach, and then adding unclustered points to the initial $K$ partial clusters in a greedy manner to form a complete clustering. To extract these seeds, we sequentially find seeds of decreasing density. The intuitive idea of sequentially finding disjoint clusters of decreasing density has been employed in various work [11, 12]. However, to the best of our knowledge, our algorithm for finding $K$ seeds and our algorithm for expanding the seeds to form complete clusters have not previously appeared in the literature. Furthermore, we are not aware of any existing mathematical analysis of density-based clustering algorithms that give guarantees similar to those presented in this work.
@@ III Preliminaries
We say a point $x_{1}$ is $\epsilon$-connected to a point $x_{t}$ if there exists some sequence of points $x_{1},x_{2},...,x_{t}$ such that $x_{i+1}$ is distance at most $\epsilon$ from $x_{i}$ for $11\leq i\leq t-$ 1 and $\epsilon_{N_{p}}(x_{i})\leq\epsilon$ for $1\leq i\leq t$. A set of points $c$ is called $\epsilon$-connected if every pair of points in $c$ is $\epsilon$-connected.
For a given $\epsilon$, a set of points $c$ is called an $\epsilon$-cluster if it is a maximal $\epsilon$-connected set of points. Any point that is $\epsilon$-connected to a point in an $\epsilon$-cluster $c$ is included in $c$. A set of points $c$ is called a maximal cluster if it is an $\epsilon$-cluster for some $\epsilon$. For a given $\epsilon$, it is helpful to consider the graph that is formed where each node corresponds to a point in $X$, and an edge is drawn between two nodes if the corresponding points $x_{1},x_{2}$ are such that $\epsilon_{N_{p}}(x_{1})\leq\epsilon$, $\epsilon_{N_{p}}(x_{2})\leq\epsilon$, and $d(x_{1},x_{2})\leq\epsilon$. A set of points is an $\epsilon$-cluster if and only if it corresponds to a connected component in this graph.
@@
For a node $v\in V$, $c(v)$ denotes the cluster corresponding to $v$. Each node $v$ has a real number $\epsilon(v)$ associated with it, where $\epsilon(v)$ is the smallest $\epsilon$ such that $c(v)$ is a $\epsilon$-cluster. For a leaf node $v$ corresponding to the cluster $\{x\}$, $\epsilon(v)=\epsilon_{N_{p}}(x)$. A node $v_{0}$ has children $v_{1},v_{2},...,v_{i}$ if $c(v_{1}),c(v_{2}),...,c(v_{i})$ form the smallest possible partition of $c(v_{0})$ composed of maximal clusters, which is guaranteed to be unique. The root node $v_{r}$ of the resulting tree is such that $c(v_{r})=X$.
@@ IV-A Weak Separability
For a given $N_{p}$, $C$ is called weakly separable if for each $c\in C$, $c$ is $\epsilon$-connected for some $\epsilon<\min_{c^{\prime}\in C,\;c^{\prime}\neq c}\epsilon(c,c^{\prime})$.
If a clustering $C$ is not weakly separable, there is ambiguity in what the correct clustering should be in the sense that there must exist a cluster $c\in C$ such that any $\epsilon$-cluster containing $c$ must include at least one point from a distinct cluster $c^{\prime}\in C$. Figures 2 and 3 illustrate weak separability.
@@
If there are two weakly separable clusterings for $X$, it is impossible to guarantee that we recover one of them if our only criterion is to find a weakly separable clustering. Because there usually does not exist a unique weakly separable $K$-clustering even when an intuitively correct $K$-clustering exists, weak separability alone as a sufficient condition for clustering recoverability is not adequate.
@@ IV-B Local Maximum Separability
LM-separability specifies that for every local maximum $x\in X$ where there exists a $y\in c(x)$ whose density is at least as high as that of $x$, the relative separability of $x$ to any such $y$ is smaller than the relative separability of any highest density point in a cluster to another cluster. This is a condition that should hold for a clustering to unambiguous because if there exists a local maximum $x$ where the relative separability to such a point $y\in c(x)$ is higher than the relative separability of a highest density point $z\in X_{c}^{*}$ for some $c\in C$ to another $c^{\prime}\in C$ (and therefore to a local maximum in $c^{\prime}$), then in a sense, $x$ looks more like it belongs to a separate cluster from $y$ than $z$ looks like it belongs to a separate cluster from $c^{\prime}$. Figures 1, 2, and 3 illustrate LM-separability.
Figure 1: A plot showing the sparsity value $\epsilon_{3}(x)$ of each point $x\in X$ in a 14 point one-dimensional clustering $C=[c_{1},c_{2}]$ that is weakly separable and LM-separable for $N_{p}=3$, implying that recovery is guaranteed. The red points correspond to points in $X_{c}^{*}$ for $c\in C$. The pink point corresponds to the local maximum $x\in X$ such that $\max_{y\in c(x):\;\epsilon_{N_{p}}(y)\leq\epsilon_{N_{p}}(x)}A(x,y)=A^{\ell}(C)$ (i.e. the local maximum that looks the most separated from its cluster). $C$ is weakly separable because $\epsilon^{*}(c_{1})=0.9$ and $\epsilon^{*}(c_{2})=0.5$, while $\min_{c,c^{\prime}\in C}\epsilon(c,c^{\prime})=1.1$. $C$ is LM-separable because $A^{\ell}(C)=2.0$ while $\min_{c,c^{\prime}\in C}\min_{z\in X_{c}^{*}}A(z,c^{\prime})=4.4$.
Figure 2: A plot showing the sparsity value $\epsilon_{3}(x)$ of each point $x\in X$ in a 60 point one-dimensional clustering $C=[c_{1},c_{2}]$ that is weakly separable and LM-separable for $N_{p}=3$, implying that recovery is guaranteed. The orange, red, and pink points correspond to local maxima. The red points correspond to points in $X_{c}^{*}$ for $c\in C$. The pink point corresponds to the local maximum $x\in X$ such that $\max_{y\in c(x):\;\epsilon_{N_{p}}(y)\leq\epsilon_{N_{p}}(x)}A(x,y)=A^{\ell}(C)$ (i.e. the local maximum that looks the most separated from its cluster). $C$ is weakly separable because $\epsilon^{*}(c_{1})=0.74$ and $\epsilon^{*}(c_{2})=1.14$, while $\min_{c,c^{\prime}\in C}\epsilon(c,c^{\prime})=1.18$. $C$ is LM-separable because $A^{\ell}(C)=20.49$ while $\min_{c,c^{\prime}\in C}\min_{z\in X_{c}^{*}}A(z,c^{\prime})=44.27$.
Note that LM-separability does not imply weak separability. Let $N_{p}=2$, and consider the clustering $C=[\{7,8,10,13,21\},\{17,25,27\}]$. $C$ is clearly not weakly separable because the clusters overlap, but is LM-separable, as $7,8,25,27$ are the only local maxima.
@@
For a given $N_{p}$, it is in general not possible to recover the unique weakly separable, LM-separable clustering by simply finding an extension of the (partial) clustering given by the $\epsilon$-cut of the dendrogram that contains $K$ clusters (if there exists such an $\epsilon$) as proved in Lemma 2. Since DBSCAN follows this approach, it is not sufficient for finding the unique weakly separable, LM-separable clustering.
@@ V Algorithm and Proof of Theorem 1
Algorithm 1 Minimal Seed Expansion (MSE)
@@
1:  $C^{\prime}\leftarrow$ min-$A$ clustering of $X$ for $N_{p},M,D,K$
2:  $\hat{C}\leftarrow$ output of Algorithm 3 with input $X,C^{\prime},N_{p}$
@@
The min-$A$ clustering of $X$ for $N_{p},M,D,K$ is given by $C_{g}(X,N_{p},A_{min},M,D)$ where
$A_{min}=\min\{A:|C_{g}(X,N_{p},A,M,D)|=K\}.$
For parameters $N_{p},M,D,K$, the min-$A$ clustering is denoted by
$C_{m}(X,N_{p},M,D,K).$
$A_{min}\geq 1$ since $c^{*}(x,\epsilon)=\{\}$ if $\epsilon<\epsilon_{N_{p}}(x)$.
Algorithm 1 accepts $N_{p},M,D$ along with a number of clusters $K$ as input, and begins by finding the min-$A$ partial clustering for $N_{p},M,D,K$. The min-$A$ partial clustering is then passed to Algorithm 3 to produce the final output clustering.
@@
        $\text{Tried}\leftarrow\text{Tried}\cup\{x^{*}\}$
@@
  $Y\leftarrow X\setminus(\cup_{c\in C^{\prime}}c)$
@@
For a given $N_{p},$ if $C$ is LM-separable, then
$|C_{g}(X,N_{p},A^{\ell}(C),1,\infty)|=K.$
@@
We will prove that all partial clusters output by Algorithm 2 are subsets of distinct clusters in $C$. Consider the $i$th partial cluster output by Algorithm 2. The highest density point $x$ that the algorithm uses to build the $i$th cluster is clearly a local maximum, and can either be from a previously partially reconstructed cluster in $C$, or can be from a new cluster in $C$. If $x$ is from a previously partially reconstructed cluster $c^{\prime}\in C$, then the $i$th partial cluster must include some $z\in X_{c^{\prime}}^{*}$ from a previously output partial cluster by the definitions of LM-separability and $A^{\ell}(C)$, and the fact that $\epsilon_{N_{p}}(x)\geq\epsilon_{N_{p}}(z).$ This is a contradiction. This, along with the fact that $M=1\leq N_{p}$, guarantees that $x$ must be from a new cluster $c^{\prime\prime}\in C$ and $x\in X_{c^{\prime\prime}}^{*}$. The $i$th partial cluster $c^{*}(x,A^{\ell}(C)\cdot\epsilon_{N_{p}}(x))$ must only include points from $c^{\prime\prime}$ by the definition of LM-separability and $A^{\ell}(C)$. Because only a partial cluster of a new cluster in $C$ can be output in each iteration of the algorithm, a partial $K$-clustering is output. ∎
@@
Therefore, the algorithm outputs non-overlapping partial clusters such that each partial cluster chosen does not include any points from true clusters that are not yet part of a previously output partial cluster. Observe that, if no points in a cluster $c\in C$ have been added to previously output partial clusters by the time $w\in X_{c}^{*}$ is selected as the highest density unclustered point by the algorithm, then a partial cluster centered at $w$ that is a non-empty subset of $c$ will be output by the algorithm in that step since $M=1\leq N_{p}$, $A_{min}\leq A^{\ell}(C)$ and $C$ is LM-separable. Therefore, we obtain a partial clustering that contains a partial cluster centered at $w\in X_{c}^{*}$ that is a non-empty subset of $c$ for each $c\in C$. This is a min-$A$ partial clustering for $K$, so it is a partial $K$-clustering that contains a partial cluster centered at some $w\in X_{c}^{*}$ that is a non-empty subset of $c$ for each $c\in C$. Thus, the algorithm outputs a partial $K$-clustering that is extendable to $C$. ∎
@@
For a given $X,N_{p},M,D,$ the number of clusters in $C_{g}(X,N_{p},A,M,D)$ monotonically decreases as $A$ increases because of the following property. Consider some $a,a^{\prime}\geq 1$ such that $a>a^{\prime}$. At the end of the $i$th iteration of Algorithm 2, the set of remaining candidates for future $x^{*}$ where $c^{*}(x^{*},A\cdot\epsilon_{N_{p}}(x^{*}))$ will not intersect with a previously output cluster in the case when $A=a^{\prime}$ is a superset of the set of such candidates when $A=a$.
@@
Observe that for each greedy step of Algorithm 3, any point $x\in X\setminus(\cup_{c\in\hat{C}}c)$ and cluster $c\in\hat{C}$ that minimize the quantity $\epsilon^{!}(x,c)=\min_{y\in c}\max(d(x,y),\epsilon_{N_{p}}(x),\epsilon_{N_{p}}(y))$ can be selected as $(x^{*},c^{*})$. To see this, consider the set $S$ of unclustered points that are closest to some cluster in terms of $\epsilon$-distance. Denote this minimum $\epsilon$-distance by $\epsilon^{\prime}$. Clearly, $S$ must include an unclustered point that contains a clustered point that it is $\epsilon^{\prime}$-connected to in its $\epsilon^{\prime}$-ball. Thus, at each greedy step of Algorithm 3, we can pick $(x,c)$ that minimizes $\epsilon^{!}(x,y)$.
@@
For each greedy step of Algorithm 3, we simply set $(x^{*},c^{*})$ equal to the $(x,c)$ that minimizes $\arg\min_{c\in\hat{C}}\epsilon^{!}(x,c)$ in $O(|X|)$ time by checking $(\epsilon^{!}(x)$, $c^{!}(x))$ for all $x\in X\setminus(\cup_{c\in\hat{C}}c)$. After assigning $x^{*}$ to $c^{*}$, we update $\epsilon^{!}(x)$ for every remaining unclustered point $x$ by setting
$\epsilon^{!}(x)\leftarrow\min(\epsilon^{!}(x),\;\max(d(x,x^{*}),\;\epsilon_{N_{p}}(x),\;\epsilon_{N_{p}}(x^{*}))),$
and setting $c^{!}(x)\leftarrow c^{*}$ if the value of $\epsilon^{!}(x)$ is changed. Each of these greedy steps takes $O(|X|)$ time and there are at most $O(|X|)$ greedy steps. This stage of the algorithm therefore runs in $O(|X|^{2})$ time. ∎